. The Trio Index is comprised of three stocks, Eins, Zwei, and Tri. Their current prices are as follows:

Stock

Eins

Zwei

Tri

Price at time t

$12

$22

$44

a.    Between now and the next time period, the stock prices of Eins and Zwei increase 10 percent while Tri increases 20 percent. What is the percentage change in the price-weighted Trio Index?
Answer #1

Since it is a price weighted index,

Weights of the respective share would be-

Eins=( 12/(12+22+44))= (12/78)= 15.38%

Zwei = (22/78)= 28.20%

Tri =( 44/78)= 56.42%

Show the changes in their price would be accordingly weighted-

Total change in the index should be=

(10*15.38%)+(10*28.20%)+(20%*56.42%)

=(1.538+2.82+11.28)

=15.638%

required change in index would be 15.638%